A grasshopper can jump maximum distance of 1.6m.it spends negligible time on ground.How far can it go in 10 seconds

Grade:11

2 Answers

Arun
6 years ago
Dear student
 
grasshopper jumps maximum distance when Ф = 45 deg

Horizontal component of velocity = u /√2 

R = U² / g         u = √gR  = √10*1.6  = 4 m/sec

Distance in 10 sec:  4/√2 * 10  = 20√2 m
 
Regards
Arun (askIITians forum expert)
Vivek Singh
4 years ago
A body is projected at time ( t=0) from a certain point on a planet's surface with a certain velocity  at a certain angle with the planet's surface ( assumed horizontal ) . the horizontal and vertical displacement x and y  ( in meter ) respectively vary with time t in second  as  , x= 10t  and y=10 t- tsquare . Then the maximum height attained by the body is